OpenStudy (anonymous):

If a simple resistor with a value of 1000 ohms is placed across the terminals of a 10 volt batter, the resistor would dra how many amps?

OpenStudy (anonymous):

sorry draw

OpenStudy (anonymous):

lol physics I = V/R I = 10/1000 I = 0.01 amphere

OpenStudy (anonymous):

thank you, so let me double check something with you if r1 = 25 and r2 = 75 when a 50 volt batter is placed across the terminals at rt the total current drawn from the batter is 5.0 E -1? or am I wrong?

OpenStudy (anonymous):

the resistors are connected in series or parallel

OpenStudy (anonymous):

mathematicalpro's question is important. If they are connected in series - meaning that there is only one path along the wires that current can take, then the resistors just add together and you can use the formula he gave you earlier replacing R with R1+R2. If they are connected so that the current can go through just one and get bact to the other side of the terminal, (called a parallel connection) then you have to replace R in his original equation with\[R =\frac {R _{1} R _{2}}{R _{1} + R_{2}}\]